UT Sep 11 2022 - Night #10 of Restart




Night Info

  • Sunset 18:32
  • Nautical Twilight Ends 19:26
  • Astro Twilight Ends 19:56
  • Astro Twilight Begins 04:37
  • Nautical Twilight Begins 05:07
  • Sunrise 06:01
  • Moonrise 19:03
  • Monnset 07:19
  • Moon illumination ~99%

Staff

  • OSA: J.Williams
  • ISp: J.Schuch
  • Mtn Manager: T.Castro
  • ISA/SO: J.Power, B.Rothberg, R.T.Gatto
  • Technical Expert: J.Hill
  • AO: G.Taylor

Night Plan

ALTA prediction for Tonight

20220910_see_evol_time_Height_20_20000.png

Starting Configuration LUCIs

Telescope Wakeup Continued

  • Conifrmatory Field Aberration (2 hrs) - Critical (1hour) - Top Priority
    • Analysis Completed
    • RFBG
      From 45 Astigmatism Measurement
      On-Sky Arcseconds
      -----------------------------
      --------------
      Tip = 4.3", Delta IE = -4.3"
      Tilt = 24.9", Delta CA = -24.9"

      rfbg
      From 0 Astigmatism Measurement
      On-Sky Arcseconds
      -----------------------------

      --------------
      Tip = -3.3", Delta IE = 3.3"
      Tilt = 27.3", Delta CA = -27.3"

      I suggest Tip = 0, Tilt = 25

      LFBG

      => Field Data Results
      Measured bi-nodal astigmatism

      lfbg
      From 45 Astigmatism Measurement
      On-Sky Arcseconds
      -----------------------------

      --------------
      Tip = 19.4", Delta IE = -19.4"
      Tilt = 16.9", Delta CA = -16.9"


      lfbg
      From 0 Astigmatism Measurement
      On-Sky Arcseconds
      -------------------------------------------
      Tip = 7.4", Delta IE = -7.4"
      Tilt = 18.0", Delta CA = -18.0”

      I suggest Tip =15, Tilt = 17 (or non at all…might be in the noise)
      • add the values to the global offsets in the M2 collimation files: /lbt/tcs/current/psf/configuration/PSF
      • SXSMLUCIFERColllimation.dat, then propagate to all the others with .csh script.
    • Can be done in parallel (data collection, analysis in serial)
    • need both AGW1 and AGW2 on the telescope and checked to execute - DOES NOT NEED LUCIs - Just AGW 1 and AGW 2
    • 1.2" or better seeing (1" or better ideal)
    • Target Fields:
      • BS9177 18:39:33.7 +49:05:38.0
      • StoneP1 22:37:45.0 -01:13:40.45
      • BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0

LUCI Checkout

  • Collect transform data - (2hrs) - confirmatory - Requires LUCIs - not complete yet
    • 1.2" or better seeing needed
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
    • Note the LUCI frame numbers associated with each offset in IDL script
    • Can be done wth LUCI SW in crippled state (does not need flex comp, remember to load EmptyQueue.xml)
  • If Transform data not needed run these - (L1 and L2 versions exist for all scripts)
    • LUCI Scripts:
      • LUCI_2022_Restart_V1.80.xml - all scripts (except LS Rot ZP)
      • Rot_ZP_V2 script
      • LUCI_Phot_ZP_OT_V6.xm - LUCI photometric standards (fixes AGW patrol problem)
    • G210 N1.8 LS Testing (Flexure and wavelength are completed) - medium/high priority
      • Choose one - scripts are designed for 300sec x 12 dithers - only need to run 4 dithers (AB pattern and make sure dithers move ok)
        • NGC 6052 - twilight-22hr
        • Arp 230 - 2nd half
    • Imaging check (1hr) - medium priority - if short on-time skip
      • ANY seeing OK
      • dither in-field
        • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am - only need to run 4 dithers
      • dither out-of-field (run one)
        • M52 - JHK, dither 5' out of field - visible all night
        • M40 - Broad filters, few narrow filters - visible all night

AO

  • Medium Priority
  • If low on time - switch to LBCs to get the TMS testing done - can do AO/ESM testing during PETS split night or UT Sep 21 or Sep 22
  • Scripts:
    • AO: /home/lbto/Service_Observing/Restart_2022/LUCI/AO/AO_Testing_V94.xml
    • ESM: /home/lbto/Service_Observing/Restart_2022/LUCI/ESM/ESM_V44.xml
  • Full AO
    • Bin 3 - pick a star from the OT that matches Bin and visibility - check to make sure GS is out of patrol field on off dithers
    • Bin 4 - pick a star from the OT that matches Bin and visibility - check to make sure GS is out of patrol field on off dithers
  • ESM
    • run as mixed mode so SL data can be taken for comparison, mixed mode scripts already exported
    • mono versions exported in case of L2 problems
    • Mixed mode scripts:
      • Bin 1 - M52 R=10.7 (11.9) AO Ref star - up all night - KHJ - can run just K if low on time
        • /home/lbto/Service_Observing/Restart_2022/LUCI/ESM/M52
      • Bin 2 - M34 R=12.4 (13.6) AO Ref star - up 22:30-twilight - KHJ - can run just K if low on time
        • /home/lbto/Service_Observing/Restart_2022/LUCI/ESM/M34
      • Bin 4 - M52 Bin 4 (forced to R=16.51) AO Ref star - up all night - KHJ - can run just K if low on time
        • /home/lbto/Service_Observing/Restart_2022/LUCI/ESM/M52

LBC Chcekout (if time - and/or AO not ready)

  • Reconfigure to LBCs by 03:21 local time, unless Field Aberration or AGW Transforms are still TBD
  • TMS Testing - 1 hours - High Priority - need to make sure TMS operates on-sky OK
    • Any field in OT that has plenty of stars
  • Superfocus sequence - only if time
    • Requires 1.2" or better stable seeing
  • Photometric Standards - backup - OR - use with TMS testing if can
    • any seeing, must be clear
  • Skyflats if clear

Night Log

Googly Docs link for Night Log

-- BarryRothberg - 10 Sep 2022
Topic revision: r3 - 11 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback